OpenStudy (anonymous):

use the square root property to solve the quadratic equation 7x^2-35=0

OpenStudy (anonymous):

7x^2 = 35 x^2 = 5 x = +/- (5)^1/2

OpenStudy (anonymous):

the answer choices are a. x= +or - sqrt of 5 b. x= sqrt of 5 c. x= + or - sqrt 35 / 7 d. x= + or - sqrt 245 / 7